OpenStudy (anonymous):

find the x-and y-intercepts of the equation 2x-4y=-8

OpenStudy (anonymous):

To find x-intercept let y = 0 and solve for x,

OpenStudy (anonymous):

2x - 4(0) = -8

OpenStudy (anonymous):

2x= -8

OpenStudy (anonymous):

x= -4

OpenStudy (anonymous):

Therefore, the x-intercept is (-4,0)

OpenStudy (anonymous):

To find y-intercept let x = 0 and solve for y

OpenStudy (anonymous):

2(0) -4y =-8

OpenStudy (anonymous):

-4y = -8

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

Therefore, the y-intercept is (0,2)

OpenStudy (anonymous):

how to graph it?

OpenStudy (anonymous):

The co-ordinates are (0,2) (-4,0)

OpenStudy (anonymous):

So

OpenStudy (anonymous):

equation----------------> 2x-4y=-8

OpenStudy (anonymous):

It is given

OpenStudy (anonymous):